One music download store charges a monthly fee of $10 plus $1 per song downloaded. Another music download store charges a monthly fee of $30 for all the songs you want to download.
a. How many songs would you have to download from the first store for the cost to be the same as the second store?
b. If you only download 15 songs per month, from which download store would you buy your music?


Please help me write the first line as an equation, I have tried and gotten s-10 = 30s.

Answer by richwmiller(17219)

10+s=30
a) s=20 songs
b) the first 10+s=10+15=$25